# TFLAG: Towards Practical APT Detection via Deviation-Aware Learning on Temporal Provenance Graph

### v1

@@ Abstract
Advanced Persistent Threat (APT) have grown increasingly complex and concealed, posing formidable challenges to existing Intrusion Detection Systems in identifying and mitigating these attacks. Recent studies have incorporated graph learning techniques to extract detailed information from provenance graphs, enabling the detection of attacks with greater granularity. Nevertheless, existing studies have largely overlooked the continuous yet subtle temporal variations in the structure of provenance graphs, which may correspond to surreptitious perturbation anomalies in ongoing APT attacks. Therefore, we introduce TFLAG, an advanced anomaly detection framework that for the first time integrates the structural dynamic extraction capabilities of temporal graph model with the anomaly delineation abilities of deviation networks to pinpoint covert attack activities in provenance graphs. This self-supervised integration framework leverages the graph model to extract neighbor interaction data under continuous temporal changes from historical benign behaviors within provenance graphs, while simultaneously utilizing deviation networks to accurately distinguish authentic attack activities from false positive deviations due to unexpected subtle perturbations. The experimental results indicate that, through a comprehensive design that utilizes both attribute and temporal information, it can accurately identify the time windows associated with APT attack behaviors without prior knowledge (e.g., labeled data samples), demonstrating superior accuracy compared to current state-of-the-art methods in differentiating between attack events and system false positive events.
@@ I Introduction
Recent scholarly endeavors have significantly advanced the transformation of audit logs [3] into provenance graphs, effectively redefining the challenge of APT detection as a multi-domain integration issue [4, 5, 6, 7, 8, 9]. The central issue in contemporary research is the effective utilization of extensive node and structural attribute information in provenance graphs to devise models capable of detecting APT attacks. The majority of approaches exploit the distribution of entities and events, as well as neighboring relationships within a provenance graph, to detect anomalous attacks [4, 5, 6, 10, 11]. In detailed attack scenarios, these techniques extract structural information through nodal topological relationships, semantic properties, and the extensive distribution of edges, thereby facilitating the development of models capable of identifying aberrant structures. Given that APT attack often mask their presence by reducing occurrence rates and blending seamlessly into normal system operations, reliance solely on structural data may not result in accurate detection.
@@ II Related Work
We have investigated the detection of Advanced Persistent Threats in provenance graphs through the application of graph neural networks. Consequently, our focus has shifted toward a review of the relevant literature that discusses these aspects.
@@ II-A APT Detection on Static Provenance Graph
To mitigate this limitation, recent advancements have incorporated sophisticated graph learning techniques into APT anomaly detection [4, 5, 10, 11, 12, 13], offering a more versatile and effective solution. For instance, early applications such as ThreaTrace [5] have utilized the GraphSAGE [18] model for node-level anomaly detection, focusing primarily on understanding neighborhood structures and identifying anomalies through these behavioral patterns. R-CAID [4] has introduced root node embedding methods, connecting each graph node back to its foundational root cause and employing graph attention mechanisms for dynamic classification. This method proficiently captures the structural nuances of trace graphs while minimizing disturbances from unrelated nodes. However, these systems generally overlook the impact of temporal dynamics on graph structures. MAGIC [10] employs a graph autoencoder (GraphMAE [19]) to learn the graph’s characteristics by masking benign nodes, with joint oversight from masked feature reconstruction and sample-based structural reconstruction to detect anomalies through node outliers. Nevertheless, MAGIC does not consider potential alterations in graph structure due to extended network traffic over time. While Flash [11] acknowledges spatial and temporal contextual semantics in data modeling, its semantic analysis is completed ahead of model training, thus leading to temporally discrete data. This could result in overlooked significant APT attack features demanding continuous temporal relevance, leading to potentially incomplete data capture during model training.
@@ II-B APT Detection on Dynamic Provenance Graph
Recent studies have integrated continuous time into traceability-based intrusion detection systems. Among these, Unicorn [12] epitomizes an initiative by introducing flow graphs that transform provenance graphs into incrementally updatable sketches, capturing distinct temporal snapshots of the system. This approach enables Unicorn to monitor current vectors and compare them against historical sketches of typical behavior to detect anomalies. However, Unicorn faces challenges in detecting rare, subtle anomalies obscured within numerous sketches of normal activities. To extract historical correlation information more efficiently from provenance graphs, TBDetector [20] employs a Transformer-based method to derive feature sequences from the transformed provenance graphs. The anomaly score of the sequence is calculated using both similarity scores and isolation scores. While this approach resembles Unicorn’s method of converting data into flow graphs, it similarly lacks effective use of temporal attributes and interpretability of results. In contrast, JBEIL [21] employs an inductive learning strategy using sequence diagrams. It processes log attributes through graph mapping to delineate the structure of these diagrams, focusing on the static phases of lateral movements in APT attacks. Regrettably, detection modules tailored for specific attack phases may not be as effective when applied to complex, real-world scenarios. Meanwhile, Kairos [13] harnesses the TGN [22] model, lauded for its effectiveness in dynamic graphs, to advance provenance graph anomaly detection. It develops dynamic node features through a memory module that predicts edge categories and continually refines the model in conventional settings, while it identifies anomalies using reconstructed edge information. Nonetheless, Kairos is hindered by suboptimal optimization for APT data characteristics, thus heightening the potential for misclassifying benign edges as anomalous.
@@ III-A Temporal Provenance Graph
In real scenarios, data collection tools used in system monitoring generate substantial amounts of traffic data rapidly, wherein APT attacks are frequently hidden using sophisticated methods. The abundant information inherent in continuously occurring events challenges the effectiveness of coarse-grained, discrete snapshot-based data analysis. The system must analyze numerous time-segmented windows to differentiate between anomalous and normal labels for supervised training. Moreover, it struggles to identify the subtle anomaly patterns of APT attacks amidst ongoing temporal variations. To address these issues, TFLAG employs a self-supervised approach that eschews reliance on labeled anomalies. It integrates features from both source and target nodes at each timestep within the temporal flow $e(t_{i})$ of graph $G$ , to predict the type of each edge connection. Concurrently, it logs accurate prediction scores, refining the model iteratively to minimize reconstruction errors as new data emerges. Furthermore, TFLAG uses these accurate scores to train a function that maps anomaly scores, thereby effectively distinguishing between normal and attack events.
@@ III-B Motivation Example
We conduct a comprehensive analysis of the aforementioned scenario and address the issues identified in the previous studies as referenced in Section II. In the preprocessing phase, TFLAG employs hierarchical hashing technology to extract nodal attribute characteristics, which minimizes storage requirements and obviates the need for pre-training. By effectively leveraging the advantages of the TGAT model [23], TFLAG dynamically captures information about sequential neighbor events occurring around nodes and aggregates even minor perturbations within normal behavior into features. To more accurately distinguish anomalies from normal events and reduce false alarms, TFLAG not only processes sequential events but also stores the anomaly scores of aggregated features in a memory module for further fine-grained differentiation. As it reconstructs node features, TFLAG dynamically forms a deviation network from the anomaly scores of events over time, thereby increasingly distancing the attack behaviors from the network structures of normal events. This method not only considers the impact of temporal information on APT attack detection but also enhances the precise identification of attack events, which differs from previous studies that only considered structural information or utilized discrete time-based approaches. Lastly, by introducing a window-level strategy for aggregating attack events, it enables reconstruction of complete attack scenarios.
@@ V TFLAG Framework
In Module (c), reconstruction loss and deviation loss are sequentially integrated within a specific time window, enabling TFLAG to readily identify edges with the highest abnormal values during each anomalous period. By combining the reconstruction loss anomaly score, deviation loss, and node rarity, we accurately pinpoint the window with the most significant outlier (e.g., window 2 in Module (c)). Given the continuous nature of attack events and the concurrency of attack nodes, TFLAG effectively identifies all abnormal windows throughout the attack sequence.
@@ V-A Provenance Graph Construction
Fig.3 illustrates node features using one-hot encoding to depict the subject types at either end of an event. Traditional approaches to representing rich system process attributes typically result in vectors that are highly cardinal and sparse, reducing training efficiency. To overcome these challenges, we apply hierarchical hashing techniques to reduce vector sparsity by mapping high-cardinality categorical features into a fixed-dimensional vector space. The hashing function translates input features into hash values, which are then used to embed features into a fixed-dimensional vector based on these values. Moreover, hierarchical hashing involves a multi-layered approach, where each layer addresses different types of edge features, leading to a nuanced representation of the final edge attributes. This approach distinguishes subtle differences across feature categories and boosts the model’s responsiveness to variations in edge attributes through multi-level hash mapping.
@@ V-B Graph Encoder
Following the $k$-th layer, as depicted in Fig.2, the final embedding representation $\mathbb{z}_{i}(t)=\mathbf{h}_{i}^{(K)}(t)$ for nodes at both ends of an event can be reliably obtained. In the graph encoding process, the aforementioned framework enhances dynamic graph modeling capabilities through the utilization of temporal encoding and deep feature aggregation. It effectively captures temporal dependencies between nodes and associates events that are distantly separated in time.
@@ V-C Graph Decoder
Section V-B addresses the issue of encoding completeness, but recent research highlights another challenge: discerning whether subtle anomalous changes signify actual APT attacks or mere sporadic system anomalies. This issue contributes to a substantial false positive rate in many anomaly-based detection methods. For the first time, TFLAG proposes a novel approach by integrating event type prediction with deviation in network anomaly score distribution to address this problem, introducing an innovative decoder structure.
@@
In the provenance graph, the anomaly scores $A_{i}(t)$ of normal events usually display a consistent distribution [30], in contrast to the scores of attack events, which diverge markedly from this pattern. Many systems often rely on type errors or excessively high anomaly scores to identify attacks, a method that proves inadequate for detecting subtle APT attacks. This novel decoder conducts a thorough analysis of real-time events by dynamically updating the distribution of anomaly scores for normal events over time, effectively distinguishing them from anomalous events. Moreover, the decoder adjusts the overall anomaly score distribution periodically to ensure the accuracy of recent anomaly evaluations, thereby preventing outdated scores from skewing these assessments and leading to false positives. Every anomaly score recorded is timestamped and archived in the database at time $t$ as depicted in Module (b). By integrating edge type prediction with anomaly score analysis, the decoder markedly enhances the precision in detecting APT attacks while preserving high accuracy for normal events.
@@ V-D Learning Procedure
Deviation networks significantly enhance anomaly detection across various domains. However, in dynamic graphs, the deviation loss is significantly influenced by temporal factors. Over time, different events exhibit distinct abnormal variations, and using a standard normal distribution as a reference score complicates handling diverse scenarios. The model relies on an adjustable repository to generate reference scores, assuming that the distribution of anomaly scores follows a Gaussian distribution—an effective approach for various datasets [33]. Additionally, incorporating time labels introduces an attenuation effect, thereby differentially weighting statistical samples across varying time spans. With these timestamped reference scores, deviations are calculated as follows:
@@
in this formulation, $\mathbf{x}(t)$ represents the joint characteristics of the edge at time $t$, while $\mu_{R}(t)$ and $\sigma_{R}(t)$ denote the mean and standard deviation of the reference scores at that time, respectively.
@@
$+y\cdot\operatorname{sigmoid}(a-\operatorname{dev}(\mathbf{x},t)).$
@@ V-E Anomaly Detection
Anomaly Window Detection. Previous studies employing node frequency-based methods for measuring rarity were overly simplistic, generally focusing solely on the frequency of node occurrences in the data or the path’s branch count [34, 35]. This approach overlooks fundamental anomaly analysis and the influence of temporal structure. Although the state-of-the-art method [13] takes both factors into consideration, its practical application faces challenges: the reconstruction error does not effectively serve as the primary criterion for identifying abnormal time windows. Moreover, when assessing node rarity, this method relies on prior knowledge to filter out noisy nodes, which contradicts the initial goal of self-supervision in the system’s design.
The anomaly detection system of TFLAG is strategically engineered based on the characteristic behaviors of APT attacks. Exploiting TFLAG’s capability to integrate both reconstruction loss and deviation loss of subtle perturbation attacks outside the average distribution, the system significantly curtails false positives and enhances the accuracy of true attack detection, all without necessitating prior knowledge of the system’s standard operations. The system evaluates the anomaly level across an entire time window by pinpointing the most anomalous events within that period, thereby determining its potential as the focal point for comprehensive attack behavior. The operational steps include: Initial Anomaly Ranking. Time windows are prioritized based on the events that exhibit the highest anomaly reconstruction loss, resulting in a sequence of windows ranked in descending order of reconstruction loss. Refining Anomaly Detection. Following initial sorting, the top 10% of windows undergo further refinement using the deviation loss of anomaly scores. An edge characterized by high reconstruction loss yet average deviation loss is likely indicative of normal behavior that diverges from typical system activity patterns. Frequency-Based Filtering. An integral part of attack filtering involves analyzing event frequency to pinpoint the central anomalous window. APT attacks are typically marked by elusive, low-frequency behaviors, contrasting with high node frequencies, which usually signify the routine operations of specific system processes.
@@
If the specified conditions are met, the two windows are considered related to the attack event, and we proceed to rank the edges within the window by their reconstruction losses. Specifically, we extract the top 10% of edges with the highest reconstruction losses and remove the associated rare nodes. To further analyze the connectivity of edge nodes within the window, we utilize Jaccard similarity [36] to measure node similarity. This metric computes the similarity between two nodes by comparing the ratio of their intersection to their union. This computed value is then compared with the values from the last two windows. If this comparison is inconclusive, we establish a feature value as the baseline for anomaly detection.
@@
The terms $SX_{T_{i}}$ and $SX_{T_{i+1}}$ represent the sets of low-frequency nodes filtered from two adjacent time windows, respectively. The Jaccard similarity value, which ranges from 0 to 1, quantifies the similarity between these node sets; a value approaching 1 indicates greater similarity.
By calculating the Jaccard similarity between anomalous low-frequency nodes in adjacent windows, we quantitatively assess their overlap, thereby facilitating more precise identification and association of anomalous events. A higher similarity score between two windows signifies a stronger correlation, implying that they are part of the same attack event. The anomaly detection system of TFLAG not only obviates the need for prior knowledge of noise nodes but also flexibly capitalizes on the low-frequency, stealthy characteristics of APT attacks, thereby increasing the model’s event detection accuracy and diminishing false positives.

### v2

# Towards Covert APT Detection via Deviation-Aware Learning on Temporal Provenance Graph

@@ Abstract
Advanced Persistent Threat (APT) have grown increasingly complex and concealed, posing formidable challenges to existing intrusion detection systems in identifying and mitigating these attacks. Recent studies have incorporated graph learning techniques to extract detailed information from provenance graphs, enabling the detection of attacks with greater granularity. Nevertheless, existing studies have largely overlooked the continuous yet subtle temporal variations in the structure of provenance graphs, which may correspond to surreptitious perturbation anomalies in ongoing APT attacks. Therefore, we introduce TFLAG, an advanced anomaly detection framework that for the first time integrates the structural dynamic extraction capabilities of temporal graph model with the anomaly delineation abilities of deviation networks to pinpoint covert attack activities in provenance graphs. This self-supervised integration framework leverages the graph model to extract neighbor interaction data under continuous temporal changes from historical benign behaviors within provenance graphs, while simultaneously utilizing deviation networks to accurately distinguish authentic attack activities from false positive deviations due to unexpected subtle perturbations. The experimental results indicate that, through a comprehensive design that utilizes both attribute and temporal information, it can accurately identify the time windows associated with APT attack behaviors without prior knowledge (e.g., labeled data samples), demonstrating superior accuracy compared to current state-of-the-art methods in differentiating between attack events and system false positive events. Our code and datasets are available at https://github.com/wangkai-tech23/TFLAG.
@@ I Introduction
Recent scholarly endeavors have significantly advanced the transformation of audit logs [3] into provenance graphs, effectively redefining the challenge of APT detection as a multi-domain integration issue [4, 5, 6, 7, 8, 9]. The central issue in contemporary research is the effective utilization of extensive node and structural attribute information in provenance graphs to devise models capable of detecting APT attacks. The majority of approaches exploit the distribution of entities and events, as well as neighboring relationships within a provenance graph, to detect anomalous attacks [4, 5, 6, 10, 11]. In detailed attack scenarios, these techniques extract structural information through nodal topological relationships, semantic properties, and the extensive distribution of edges, thereby facilitating the development of models capable of identifying aberrant structures. Given that APT attacks often mask their presence by reducing occurrence rates and blending seamlessly into normal system operations, reliance solely on structural data may not result in accurate detection.
@@ II Related Work
We have investigated the detection of APT attacks in provenance graphs through the application of graph neural networks. Consequently, our focus has shifted toward a review of the relevant literature that discusses these aspects.
@@ II-A APT Detection on Static Provenance Graph
To mitigate this limitation, recent advancements have incorporated sophisticated graph learning techniques into APT anomaly detection [4, 5, 10, 11, 12, 13], offering a more versatile and effective solution. For instance, early applications such as ThreaTrace [5] have utilized the GraphSAGE [18] model for node-level anomaly detection, focusing primarily on understanding neighborhood structures and identifying anomalies through these behavioral patterns. R-CAID [4] has introduced root node embedding methods, connecting each graph node back to its foundational root cause and employing graph attention mechanisms for dynamic classification. This method proficiently captures the structural nuances of trace graphs while minimizing disturbances from unrelated nodes. However, these systems generally overlook the impact of temporal dynamics on graph structures. MAGIC [10] employs a graph autoencoder (GraphMAE [19]) to learn the graph’s characteristics by masking benign nodes, with joint oversight from masked feature reconstruction and sample-based structural reconstruction to detect anomalies through node outliers. Nevertheless, MAGIC does not consider potential alterations in graph structure due to extended network traffic over time. Although Flash [11] takes into account the spatial and temporal contextual semantics in data modeling, its semantic analysis is performed prior to model training, which results in temporally discrete data. This could result in overlooked significant APT attack features demanding continuous temporal relevance, leading to potentially incomplete data capture during model training.
@@ II-B APT Detection on Dynamic Provenance Graph
Recent studies have integrated continuous time into traceability-based intrusion detection systems. Among these, Unicorn [12] epitomizes an initiative by introducing flow graphs that transform provenance graphs into incrementally updatable sketches, capturing distinct temporal snapshots of the system. This approach enables Unicorn to monitor current vectors and compare them against historical sketches of typical behavior to detect anomalies. However, Unicorn faces challenges in detecting rare, subtle anomalies obscured within numerous sketches of normal activities. To extract historical correlation information more efficiently from provenance graphs, TBDetector [20] employs a Transformer-based method to derive feature sequences from the transformed provenance graphs. The anomaly score of the sequence is calculated using both similarity scores and isolation scores. While this approach resembles Unicorn’s method of converting data into flow graphs, it similarly lacks effective use of temporal attributes and interpretability of results. In contrast, JBEIL [21] employs an inductive learning strategy using sequence diagrams. It processes log attributes through graph mapping to delineate the structure of these diagrams, focusing on the static phases of lateral movements in APT attacks. Regrettably, detection modules tailored for specific attack phases may not be as effective when applied to complex, real-world scenarios. Meanwhile, Kairos [13] harnesses the TGN [22] model, lauded for its effectiveness in dynamic graphs, to advance provenance graph anomaly detection. It aggregates dynamic node features through a memory module while identifying anomalies using reconstructed errors. However, Kairos is hindered by its inability to effectively distinguish specific low-frequency APT attacks hidden within a large volume of benign data and occasional system behaviors, thus increasing the likelihood of misclassifying benign edges as anomalous.
@@ III-A Temporal Provenance Graph
In real scenarios, data collection tools used in system monitoring generate substantial amounts of traffic data rapidly, wherein APT attacks are frequently hidden using sophisticated methods. The abundant information inherent in continuously occurring events challenges the effectiveness of coarse-grained, discrete snapshot-based data analysis. The system must analyze numerous time-segmented windows to differentiate between anomalous and normal labels for supervised training. Moreover, it struggles to identify the subtle anomaly patterns of APT attacks amidst ongoing temporal variations. To address these issues, TFLAG employs a self-supervised approach that eschews reliance on labeled anomalies. It integrates features from both source and target nodes at each timestep within the temporal flow $e(t_{i})$ of graph $G$ , to predict the type of each edge connection. Concurrently, it logs accurate prediction labels to build a deviation network, iteratively optimizing the model as new data emerges to minimize reconstruction errors, thereby effectively distinguishing between normal and attack events in the future.
@@ III-B Motivation Example
We conduct a comprehensive analysis of the aforementioned scenario and address the issues identified in the previous studies as referenced in Section II. By effectively leveraging the advantages of the TGAT model [23], TFLAG dynamically captures information about sequential neighbor events occurring around nodes and aggregates even minor perturbations within normal behavior into features. In contrast to TGN, which retains historical node feature information through a memory module, TGAT aggregates the temporal features of neighboring nodes and the node in focus at the current moment. This approach provides an advantage in capturing sudden interactions between network nodes (Nginx connect) and file access (vUgefal read) events. During the initial stages of an attack, this scenario generates a substantial volume of IP interactions, encompassing both normal system behaviors and the actions of the attacker. Consequently, if reconstruction is the only factor considered in detection, it becomes challenging to differentiate between system false positives and real attacks. To more accurately distinguish anomalies from normal events and reduce false alarms, TFLAG not only processes sequential events but also stores the anomaly scores of aggregated features in a memory module for further fine-grained differentiation. As it reconstructs node features, TFLAG dynamically forms a deviation network from the anomaly scores of events over time, thereby increasingly distancing the attack behaviors from the network structures of normal events. This method not only considers the impact of temporal information on APT attack detection but also enhances the precise identification of attack events.
@@ V TFLAG Framework
In Module (c), the reconstruction loss and deviation loss are sequentially integrated within a specific time window, allowing TFLAG to accurately identify edges with the highest abnormal values during each anomalous period. As events deviating from normal system behavior can also result in elevated anomaly scores, false attack windows may appear within the candidate time window queue. To effectively discern the true attack windows, we combine the reconstruction loss anomaly score, deviation loss, and node rarity, utilizing these integrated metrics to identify authentic attack windows marked by significant outliers (e.g., window 2 in Module (c)). Building on this, TFLAG recursively evaluates adjacent anomalous windows preceding and succeeding the central attack window to ascertain whether they represent false positives, ultimately enabling a comprehensive tracing of the complete attack sequence.
@@ V-A Provenance Graph Construction
Fig. 3 illustrates node features using one-hot encoding to depict the subject types at either end of an event. Traditional approaches to representing rich system process attributes typically result in vectors that are highly cardinal and sparse, reducing training efficiency. To overcome these challenges, we apply hierarchical hashing techniques [14] to reduce vector sparsity by mapping high-cardinality categorical features into a fixed-dimensional vector space. The hashing function translates input features into hash values, which are then used to embed features into a fixed-dimensional vector based on these values.
@@ V-B Graph Encoder
TGAT addresses the dispersed, phase-based nature of attacks more effectively than continuous-time models like TGN. APT attacks are characterized by intermittent, abrupt events and sudden structural changes in network graphs, such as unexpected new connections or anomalous node interactions. While TGN continuously updates embeddings for long-term dependencies, TGAT’s temporal aggregation approach better captures the sporadic bursts and phase-based behaviors typical of APT campaigns, making it superior at detecting the short-term changes that continuous models may miss.
@@
Following the $k$-th layer, as shown in Module (b), the final embedding representation $\mathbb{z}_{i}(t)=\mathbf{h}_{i}^{(K)}(t)$ for nodes at both ends of an event can be reliably obtained. In the graph encoding process, the aforementioned framework enhances dynamic graph modeling capabilities through the utilization of temporal encoding and deep feature aggregation.
@@ V-C Graph Decoder
Section V-B addresses encoding completeness, but recent research reveals another challenge: distinguishing between subtle anomalous changes indicating actual APT attacks versus sporadic system anomalies. This contributes to substantial false positive rates in anomaly-based detection methods. TFLAG addresses this by integrating event type prediction with deviation analysis in network anomaly score distribution through an innovative decoder structure.
@@
In provenance graphs, normal events typically exhibit consistent anomaly score distributions $A_{i}(t)$, while attack events show significant deviations. Traditional systems relying solely on anomaly scores or reconstruction losses produce high false positive rates. TFLAG’s novel decoder generates dynamic labels by predicting event categories and constructs a deviation network that continuously records normal event anomaly score distributions during learning. This realigns deviating scores within the established normal distribution range, enhancing system robustness. The decoder periodically recalibrates the overall anomaly score distribution to maintain accuracy for recent abnormal events, preventing outdated scores from distorting evaluations. Through dynamic event prediction and dual deviation network optimization, TFLAG accurately detects APT attacks while effectively minimizing false positives.
@@ V-D Learning Procedure
Deviation networks significantly enhance anomaly detection across various domains. However, in dynamic graphs, the deviation loss is significantly influenced by temporal factors. Over time, different events exhibit distinct abnormal variations, and using a standard normal distribution as a reference score complicates handling diverse scenarios. The model relies on an adjustable repository to generate reference scores, assuming that the distribution of the initial anomaly scores follows a Gaussian distribution—an effective approach for various datasets [32]. Additionally, incorporating time labels introduces an attenuation effect, thereby differentially weighting statistical samples across varying time spans. With these timestamped reference scores, deviations are calculated as follows:
@@
in this formulation, $\mathbf{x}(t)$ represents the aggregated features of the edge at time $t$, $\mathbf{\Theta}$ denotes the weights and biases of the multilayer perceptron, $\mathbf{\phi}(\mathbf{x}(t);\Theta)$ represents the anomaly score of the event , while $\mu_{R}(t)$ and $\sigma_{R}(t)$ denote the mean and standard deviation of the reference scores at that time, respectively.
@@
$\hskip-25.00003pt+y\cdot\operatorname{sigmoid}(a-\operatorname{dev}(\mathbf{x},t)).$
@@ V-E Anomaly Detection
Anomaly Window Detection. Previous studies employing node frequency-based methods for measuring rarity were overly simplistic, generally focusing solely on the frequency of node occurrences in the data or the path’s branch count [33, 34]. This approach overlooks fundamental anomaly analysis and the influence of temporal structure. Although the state-of-the-art method [13] takes both factors into consideration, its practical application faces challenges: when connecting multiple anomaly windows, it is necessary to precompute the node frequency of event attributes and rely on prior knowledge to manually exclude event attributes that may influence the final calculation of anomaly scores, which is difficult to achieve in real-world scenarios. In contrast, the anomaly detection module of TFLAG is designed based on the low frequency and continuity of APT attacks, so it does not require the prior exclusion of normal event attributes that may influence the calculation, and can directly detect the complete attack sequence.
The operational steps include: Initial Anomaly Ranking. Time windows are prioritized based on the events that exhibit the highest anomaly reconstruction loss, resulting in a sequence of windows ranked in descending order of reconstruction loss. Refining Anomaly Detection. Following initial sorting, the top 10% of windows undergo further refinement using the deviation loss of anomaly scores. An edge characterized by high reconstruction loss yet average deviation loss is likely indicative of normal behavior that diverges from typical system activity patterns. Frequency-Based Filtering. An integral part of attack filtering involves analyzing event frequency to pinpoint the central anomalous window. APT attacks are typically marked by elusive, low-frequency behaviors, contrasting with high node frequencies, which usually signify the routine operations of specific system processes.
@@
The terms $S_{T_{i}}$ and $E_{T_{i}}$ represent the set of nodes and the set of edges within the time window. When the specified conditions are satisfied, the two windows are deemed relevant to the attack event. Subsequently, we rank the edges based on reconstruction losses and employ Jaccard similarity [35] to assess node similarity. This metric is then compared against the last two windows, and if the results remain inconclusive, a feature value is established as the baseline for anomaly detection.
@@
The terms $SX_{T_{i}}$ and $SX_{T_{i+1}}$ refer to sets of low-frequency nodes extracted from two adjacent time windows. The Jaccard similarity, which ranges from 0 to 1, quantifies the degree of similarity between these sets; a value closer to 1 signifies a stronger similarity.
By calculating the Jaccard similarity among anomalous low-frequency nodes, we can assess their overlap, enabling a more precise identification of anomalous events. A higher score implies a stronger correlation, suggesting that the nodes belong to the same attack. Therefore, the TFLAG anomaly detection system obviates the necessity for prior knowledge of noise nodes, capitalizing on the stealthy characteristics of APT attacks to enhance detection accuracy and reduce false positives.